• Main INDEX
  • Monthly INDEX
  • PREV

    User name eschulte

    Log entry time 23:57:21 on July 31, 2006

    Entry number 176597

    keyword=Swing Shift Log

    16:00 Shift starts with no beam delivery to hall. We have the 15 cm LH2
    target with 4% radiator in. Upon delivery of beam to the hall, we will
    finish the LH2 running (13M events of which we have 1.6M) and then move
    on to the LD2 radiator in running (32M events of which we start with 2M).
    This is for KIN #6 setting.

    16:49 Harry from MCC called to let us know that they think they have
    diagnosed their injector problem and have put in a "fix". They are
    testing to make sure that everything is OK and will let us know when we
    will be getting the beam back. No estimate at this time.

    17:01 MCC called to let us know they are trying to recover from gun
    problems. They are going to send a test amount of cw to hall to check
    things out.

    17:08 MCC called to ask what current we would like. We requested 14uA
    and he's hoping the machine will cooperate.

    17:10 Restarted data taking with 14uA (to give everyone a little break
    for now) on run 1938 4% radiator, 15cm LH2, 14 uA.

    17:27 Run Coordinator changes LH2 radiator in event request to 21
    Million from 31 Million.

    17:40 MCC readjusted beam positions after a series of trips.

    18:27 After 20+ solid minutes of 14uA beam the radiator temperature
    appears to be settling down to about 44.4 deg. C.

    18:34 Another HRS-L Q2 POrb alarm, which resolved itself.

    18:37 After 30+ minutes of beam (continuous) the radiator temperature
    has nearly saturated at about 45.3 deg. C.

    18:48 Increased current to 15 uA. Radiator temp. at change about 45.4
    deg. C.

    18:52 Another HRS-L Q2 POrb alarm, which resolved itself.

    20:10 Radiator temp a nice steady 48.2 deg. C after 60+ minutes of 15uA.
    This was a 3 deg. C rise for a 1uA current increase. This only
    increased the DAQ rate by 300 Hz. Raising the current to 16 uA will
    probably cause

    20:12 ROC 4 crashed. Stopped run 1943. Trying to reboot via portserver
    produces the response "no route to host". Paged Alexander who did not
    answer, so paged B. Michaels. He responded and is trying to contact the
    portserver from home.

    20:45 Called MCC to request an access to recycle ROC 4's fastbus crate
    per B. Michaels request.

    21:28 Another HRS-L Q2 POrb alarm, which resolved itself.
    21:28 Returned from hall after power cycling the ROC 4 fastbus crate.
    The crate came back to life--at least for now and we are waiting for MCC
    to restore beam. Took some pedistal runs on both arms while we were
    waiting for the beam to be restored.

    21:35 Beam restored at 15 uA. Restarted DAQ with HRS-L run number 1946

    22:56 Another HRS-L Q2 POrb alarm, which resolved itself.

    23:13 Another HRS-L Q2 POrb alarm, which resolved itself.

    23:27 Request target change to 15 cm LD2
    23:40 Finished target change.

    23:43 Reduced current to 12uA to reduce dead-time (32% before change) on
    HRS-L DAQ. Dead-time with 12uA is about 20%.

    Shift ended with 15 cm LD2 target with 4% radiator and taking data.